Na2CO3 + Ca(OH)2--> CaCO3 + 2NAOH If the chemist mixes 75 ml of 1.5M Ca(OH)2, how many grams of NaOH will be produced?
1
Expert's answer
2020-03-31T12:31:51-0400

"n=C*V"

"n" "(Ca(OH)_2)=1.5\u041c*0.075L=0.1125mol"

The chemical reaction equation shows that sodium hydroxide is formed twice as much as calcium hydroxide reacts.

"n(NaOH)=0.1125mol*2=0.225 mol"

"m=n*M"

"m(NaOH)=0.225 mol*40g\/mol=9g"

The answer is 9 grams of NaOH